OpenStudy (anonymous):

5xy-7xz-x+6=10y+5 Solve for x

OpenStudy (anonymous):

x=0 y=1/10

OpenStudy (anonymous):

I got 10y-1/-5y+7z+1

OpenStudy (hoblos):

take x as a common factor 5xy-7xz-x+6=10y+5 x(5y -7z -1) = 10y -1 x= (10y-1)/(5y -7z -1)

OpenStudy (anonymous):

my book says you can simplify that further by multiplying by -1????????

OpenStudy (hoblos):

well i cant see how would it be simplified when we multiply by -1 !!

OpenStudy (ajprincess):

when u multiply ur answer u vl get hoblos answer

OpenStudy (anonymous):

look

OpenStudy (anonymous):

@ajprincess what where you typing?

OpenStudy (ajprincess):

i dnt get y they multiply by -1 coz hoblos answer itself is a simplified 1.

OpenStudy (ajprincess):

I feel that it's better to leave the first term as positive than making t negativ by multiplying t with -1. This was wat I typed. Is that k nw @strawberry_militia